OpenStudy (anonymous):

Which has the greatest value? 5 out of 9 zero point five thirty percent three-eighths

OpenStudy (geneticrockhopper247):

3/8=.375 0.5 30%=.3 5out of 9=5/9=.555555 Which one is the greatest and that is your answer.

OpenStudy (anonymous):

.375

OpenStudy (anonymous):

@geneticrockhopper247

OpenStudy (anonymous):

No wait it's .555555

OpenStudy (geneticrockhopper247):

correct :)